# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 3.2 Hydrodynamic Process Understanding
Hydrodynamic models are mathematical representations aimed at simulating fluid movement, typically necessitating computational solutions. These models simulate water motion by solving equations derived from the application of fundamental laws of physics, encompassing principles governing the conservation of mass, momentum, and energy. Hydrodynamic models can realize the simulation of hydrological processes (such as floods and rainfall-induced landslides) in a time scale of seconds, hours, or days. Hydrodynamic models are classified into one-dimensional (1D), two-dimensional (2D), and three-dimensional (3D) models based on their spatial representation of flow dynamics. Detailed information about Physical Methods in Hydrodynamic Modeling is provided in Appendix B of Supplementary Material.
@@ 3.2.1 Hydrodynamic Modeling Meets Physics-aware Machine Learning
Table 6: PaML-based Hydrodynamic Modeling Classified by Objectives
@@
First, PDgML shows promise for real-time hydrodynamic modeling/forecasting due to its simplicity, superior performance, and computational efficiency. Data-driven models [271, 267, 268] provide a novel phenomenological approach to learning constitutive relations of hydrodynamics from data. Specifically, CNN techniques  [269, 270] are applied in 2D hydraulic/hydrodynamic models to predict water depths using outputs from existing hydraulic models. Deep convolutional generative adversarial networks (DCGANs) have been developed for real-time flow forecasting [272]. For real-time feedback control of 2D hydrodynamics, a DRL strategy is proposed, focusing on controlling fluidic pinball dynamics and tracking force extrema resulting from cylinder rotations [273]. An adaptive neural fuzzy inference system (ANFIS) is utilized for modeling 3D flows in large-scale rivers, trained with coefficients derived from the dynamics of flood flow using large-eddy simulation [274]. In addition, for a Lagrangian form of hydrodynamics, a hydrodynamic interaction graph neural network (HIGNN) is introduced for inferring and predicting particle dynamics in Stokes suspensions [275]. A second use of PaML can be found in PiML. Cai et al. [16] review physics-informed learning of flow dynamics, which integrates data and mathematical models through the use of PINNs. Specifically, PINNs of the Saint-Venant equations [27, 29] are developed for spatial-temporal scale flood forecasting. A PINN built directly from a configuration of the Saint-Venant equations [276] is created for use in real river channels, demonstrating high prediction accuracy and scientifically consistent behavior. PINNs [28, 277] are also utilized for solving 2D shallow-water equations. NSFnets [110] is developed for the incompressible Navier-Stokes equations. Most recently, PINNs (PiML) is widely used in the field of hydrodynamics [310]. However, the current generation of PINNs lacks the accuracy and efficiency of high-order CFD codes for solving hydrodynamic equations. State-of-the-art PiML methods, including advanced PINNs for addressing the challenges of vanilla PINNs, and the physics-informed neural operators in Table 2.2, can be used to alleviate the gap between PiML and hydrodynamic equation solutions. Third, for PeML, these hydrodynamic equations (such as Lagrangian forms of Navier-Stokes equations, and differential equations) are used to design the frameworks or modules of NNs, such as FGN [159], continuous convolutions [44], and differentiable modeling [13]. In addition, the hydrodynamic model–FLO 2D, coupled with a ML algorithm-scaled conjugate gradient neural network (SCG-NN) [278], is developed by embedding flood properties (such as rainfall intensity, drainage density, soil type, land cover, and others) into the input spaces of NNs. The hydrological boundaries (such as the lake boundary and downstream boundary) can be embedded into a 1D HydroDynamic model with Gated Recurrent Units (GRU-HD) [279]. However, these PeML methods still lack interpretability. The deep embedding of data-driven methods and physics-driven methods is urgently needed. Finally, PaHL is a more flexible means of solving hydrodynamic equations. For example, parallel hybrid methods (such as the differentiable hybrid approach [192], and deep learning-based shallow water equations solver [280]) are designed to capture the hydrodynamic process of the fluid. Parallel hybrid methods (such as GRU-HD [279]) are proposed to quickly and accurately simulate the water process.
@@
Model Generalizability and Transferability. See Fig. 5 (label 2.3). Despite the high accuracy of physics-based hydrodynamic methods, these models often have extended simulation durations, limiting their utility for exploratory or real-time flood predictions. A robust and efficient PaML can replace the computationally expensive parts of hydrodynamic models by improving the generalizability and transferability of PaML models. Specifically, the NN models’ transferability for predicting water depth [284] can be further enhanced through the utilization of transfer learning techniques. A transfer learning technique is also utilized to successively correct the trained DeepONet at the prediction steps [286]. A practical data-driven forecasting method based on dynamic systems theory [285] is developed to predict unprecedented water levels. For PiML, there are many derivatives of PINNs designed to enhance generalizability and transferability for solving PDEs, such as meta-learning based PINNs [125, 126, 127] and TPINN [131]. In addition, in order to improve the generalizability and transferability of PaHL methods, Jamali et al. [288] substitute the time-intensive components of physics-based numerical models with more efficient data-driven approaches, thereby enhancing their computational efficiency and achieving robust generalization capability. Kahana et al. [287] explore the geometry transferability properties of HINTS [190].
@@
Speed and Operability. See Fig. 5 (label 2.4). Real-time and operational hydrodynamic modeling (such as flood forecasting) is crucial for supporting emergency responses and reducing risk and damage [311]. Currently, there are few operational systems capable of forecasting flooding at spatial resolutions sufficient to support emergency preparedness and response efforts aimed at mitigating the impacts of floods. Combining the fast inference of ML models with the reliability of physical models, real-time and operability of hydrodynamic modeling is possible. Specifically, a real-time predictive DCGAN [272] is developed for forecasting floods. The last one-time step-ahead forecast from the DCGAN can serve as a new input for the subsequent time step-ahead forecast, which recursively forms a long lead-time forecast. Random forest [289] serves as a surrogate model for real-time flood prediction at a street scale. The surrogate model is trained using hourly water depths simulated by a 1D/2D hydrodynamic model. In addition, this PiML-based method [283] for reducing the computational load of traditional hydrodynamic methods (such as scalable hydrodynamic modeling) can be used for real-time hydrodynamic modeling. Furthermore, a hybrid framework for real-time flood modeling that merges fluid dynamics with probabilistic techniques [290] is developed to address the excessive computational requirements of high-fidelity real-time modeling. Google’s operational flood forecasting system [220], which integrates hybrid hydrological and hydrodynamic models, is established to deliver precise real-time flood alerts to agencies and the public. This forecasting system includes data validation, stage prediction, inundation modeling, and alert dissemination.
@@
Parameterization and Calibration. See Fig. 5 (label 2.5). The behaviors of hydrodynamic models depend heavily on parameters (such as river bed elevation, channel geometry, precipitation, land use land cover, etc.) that need calibration. However, traditional calibration (such as manually-calibrated process-based models) is highly inefficient and results in nonunique solutions [227]. Thus, many PaML-based parameter learning methods are proposed for hydrodynamic modeling. Specifically, a novel genetic-k-nearest neighbor hybrid algorithm [291] is proposed to mitigate the impractically high computational efforts associated with conventional calibration search techniques, while still achieving high-quality calibration results. A set of statistical-learning-based data assimilation (SLDA) methods [292] is proposed for estimating parameters in hydrodynamic models. In addition, according to the PiML community, PINNs can be used to solve parameter identification and parameter learning (data-physics-driven parameter discovery). For example, PINNs are actively trained to approximate solutions to the Navier-Stokes equations over a parameter space region, where these parameters define physical properties like domain shape and boundary conditions [293]. Furthermore, differentiable parameter learning [227] is proposed to efficiently learn a global mapping between inputs (and optionally responses) and parameters. Differentiable parameter learning contains a parameter estimation module that maps from raw input information to process-based model parameters, as well as a differentiable process-based model. The use of differentiable parameter learning leads to improved performance, enhanced physical coherence, superior generalizability, and reduced computational cost.
@@
Data Generation. See Fig. 5 (label 2.6). Data generation methods are valuable for simulating hydrodynamic processes under specific conditions. For instance, an integrated model [294] combining hydrodynamics and ML is proposed to predict water level dynamics as an indicator of compound flooding risk in a data-limited delta region. In this integrated approach, a hydrodynamic model first simulates various scenarios of compound flooding, and the outputs are then used to train the ML model. A hybrid framework [301] that integrates a hydrodynamic model, rapid flood model, and ML model is used for flood susceptibility assessments. Flood inventory data is generated by the hydrodynamic model. However, traditional methods in hydrodynamics for generating data typically involve running physical simulations or conducting experiments, which are often time-consuming. Hence, there is growing interest in generative machine learning approaches that can learn data distributions in unsupervised settings, potentially generating novel data beyond what traditional methods can produce. For example, FloodGAN [295] utilizes an image-to-image translation approach where the model learns to generate 2D inundation predictions conditioned on heterogeneous rainfall distributions, employing a minimax game between two adversarial networks. Conditional GAN [296] can also be trained to simulate fluid flow based solely on observations, without knowledge of the underlying governing equations. However, a well-known drawback of GANs is their high sample complexity and lack of diversity. Therefore, some PeML- and PiML-based GANs that can leverage prior knowledge of physics are proposed for hydrodynamic processes. For example, GAN-based models for simulating turbulent flows can be enhanced by incorporating conservation laws into the loss function [297]. Physics-informed GAN [298] is employed to learn a functional prior from historical data. In addition, some generative modeling approaches are developed to efficiently model data spaces with known invariances [299, 300].
@@
However, most hydraulic problems use traditional hydrodynamic models to generate data. With the ongoing advancements in generative models (such as GANs and diffusion models [312]) in the PaML community, the rapid generation of physical data based on generative models will gradually dominate.
@@
Uncertainty Quantification in Hydrodynamic Modeling. See Fig. 5 (label 2.7). Estimating uncertainty in hydrodynamic modeling is important for many applications. For instance, understanding, characterizing, and quantifying the impact of model and parametric uncertainties is crucial for making informed risk-based decisions in response to intense rainfall events. Monte Carlo, first-order second-moment, and metamodeling are three popular methods to estimate the uncertainty of hydrodynamic models [313]. Abbaszadeh et al. [314] explore multiple origins of uncertainties across distinct layers of hydrometeorological and hydrodynamic model simulations, as well as their intricate interactions and cascading outcomes, such as the propagation of uncertainty. They highlight how uncertainties in forecasting flooding are influenced. Specifically, recent advances in PDgML techniques have shown the benefits of using remote sensing data to correct elevation errors associated with building artifacts, flood defense structures, forests, and wetlands (such as object-based correction [302], FABDEM [303], and random forest regression [304]). In addition, Bayesian estimation/inversion methods are commonly employed to quantify and mitigate modeling uncertainties, such as employing a Markov chain Monte Carlo approach [305] or Bayesian calibration [306]. Furthermore, PiML and PeML approaches have also been gaining attention in the scientific community, as the computational burden and efficiency required for complex hydrodynamic modeling can be alleviated by these methods. For example, the physics-informed Gaussian process regression method [307] combined with the multi-level Monte Carlo for reducing the cost of estimating uncertainty is designed for real-time flood depth forecasting. In addition, a deep probabilistic generative model [308] implements a physics-based loss for uncertainty quantification, ensuring adherence to the structure imposed by PDEs. Bayesian PINNs [113] is proposed by using a Bayesian NN, which naturally encodes uncertainty. Bayesian PINNs [113] incorporate a PDE constraint to enforce the governing laws of the system as a prior for the Bayesian Network, enhancing prediction accuracy in scenarios with significant noise through physics-based regularization. Another direction is a combination of physics knowledge and ML for uncertainty quantification. Hybrid frameworks that combine ML and hydrodynamic models can identify spatio-temporal features from historical flood events, aiding in predicting spatially distributed water depth and flood inundation extent caused by fluvial and coastal drivers [309].
@@
Integrating prior physics knowledge into machine learning for uncertainty quantification in hydrodynamic modeling has the potential to better characterize uncertainty. Further research in this area is recommended to fully leverage the benefits of the PaML community. PaML models have gained popularity in recent years as an efficient tool for hydrodynamic modeling, designing these emulators and their configurations remains challenging due to the uncertainties inherent in physical models, ML models, and their hybrid methods. There have been insufficient studies to rely on PaML and convincingly demonstrate its utility and effectiveness in solving challenging hydrodynamic problems.
Figure 6: The operational flood forecasting system based on HydroPML
3.3 Proposed HydroPML-based Flood Forecasting
@@
Based on the HydroPML technology, an operational flood forecasting system is developed. As shown in Fig. 6, four key modeling components are integrated into flood warning systems: (1) data management based on parameterization and calibration, (2) a rainfall-runoff forecasting model based on PaML, (3) a flood inundation model based on hydrodynamic models, and (4) alerts to the affected population.

### v5

@@ 3.2 Hydrodynamic Process Understanding
Hydrodynamic models are mathematical models that describes fluid movement and typically require numerical methods for accurate solutions. These models simulate water motion by solving equations derived from the application of fundamental laws of physics, encompassing principles governing the conservation of mass, momentum, and energy. Hydrodynamic models can realize the simulation of hydrological processes (such as floods and rainfall-induced landslides) in a time scale of seconds, hours, or days. Hydrodynamic models are classified into one-dimensional (1D), two-dimensional (2D), and three-dimensional (3D) models based on the spatial representation of the flow. Detailed information about Physical Methods in Hydrodynamic Modeling is provided in Appendix B of SM.
@@ 3.2.1 Hydrodynamic Modeling Meets Physics-aware Machine Learning
Table 3: PaML-based hydrodynamic modeling classified by objectives.
@@
First, PDgML shows great potential for real-time hydrodynamic modeling/forecasting due to its simplicity, superior performance, and computational efficiency. Data-driven models [279, 275, 276] provide a new phenomenological approach to learning constitutive relations of hydrodynamics from data. Specifically, CNN techniques [277, 278] are applied in 2D hydraulic/hydrodynamic models to predict water depths using outputs from existing hydraulic models. Deep convolutional generative adversarial networks (DCGANs) [280] have been developed for real-time flow forecasting. For real-time feedback control of 2D hydrodynamics, a DRL-based feedback strategy [281] is proposed, focusing on controlling the 2D hydrodynamic on fluidic pinball, i.e., force extremum and tracking, from cylinders’ rotation. An adaptive neural fuzzy inference system (ANFIS) [282] is utilized for modeling 3D flows in large-scale rivers. The learnable coefficients of ANFIS are trained on 3D flood flow dynamics data using large-eddy simulation. In addition, for a Lagrangian form of hydrodynamics, a hydrodynamic interaction graph neural network (HIGNN) [283] is introduced for inferring and predicting particle dynamics in Stokes suspensions. A second use of PaML can be found in PiML. Cai et al. [16] review physics-informed learning of flow dynamics, which integrates data and mathematical models through the use of PINNs. Specifically, PINNs of the Saint-Venant equations [27, 29] are developed for spatial-temporal scale flood forecasting. A PINN built directly from a configuration of the Saint-Venant equations [284] is created for use in real river channels, demonstrating high prediction accuracy and scientifically consistent behavior. PINNs [28, 285] are also utilized for solving 2D shallow-water equations. NSFnets [107] is developed for the incompressible Navier-Stokes equations. Most recently, PINNs (PiML) is widely used in the field of hydrodynamics [318]. However, the current generation of PINNs lacks the accuracy and efficiency of high-order CFD codes for solving hydrodynamic equations. State-of-the-art PiML methods, including advanced PINNs for addressing the challenges of vanilla PINNs, and the PINOs in Supplementary Table 2, can be used to alleviate the gap between PiML and hydrodynamic equation solutions. Third, for PeML, these hydrodynamic equations (such as Lagrangian forms of Navier-Stokes equations, and differential equations) are used to design the frameworks or modules of NNs, such as FGN [157], continuous convolutions [76], and differentiable modeling [13]. In addition, the hydrodynamic model–FLO 2D, coupled with a ML algorithm-scaled conjugate gradient neural network (SCG-NN) [286], is developed by embedding flood properties (such as rainfall intensity, drainage density, soil type, land cover, and others) into the input spaces of NNs. The hydrological boundaries (such as the lake boundary and downstream boundary) can be embedded into the gated recurrent unit with a 1D HydroDynamic model (GRU-HD) [287]. However, these PeML methods still lack interpretability. The deep embedding of data-driven methods and physics-driven methods is urgently needed. Finally, PaHL is a more flexible means of solving hydrodynamic equations. For example, parallel hybrid methods (such as the differentiable hybrid approach [191], and deep learning-based shallow water equations solver [288]) are designed to capture the hydrodynamic process of the fluid. Parallel hybrid methods (such as GRU-HD [287]) are proposed to quickly and accurately simulate the water process.
@@
Model Generalizability and Transferability. See Fig. 5 (label 2.3). Despite the high accuracy of physics-based hydrodynamic methods, these models have long simulation run times and therefore are of limited use for exploratory or real-time flood predictions. A robust and efficient PaML can replace the computationally expensive parts of hydrodynamic models by improving the generalizability and transferability of PaML models. Specifically, the NN models’ transferability for predicting water depth [292] can be further enhanced through the utilization of transfer learning techniques. A transfer learning technique is also utilized to successively correct the trained DeepONet at the prediction steps [294]. A practical data-driven forecasting method based on dynamic systems theory [293] is developed to predict unprecedented water levels. For PiML, there are many derivatives of PINNs designed to enhance generalizability and transferability for solving PDEs, such as meta-learning based PINNs [130, 131, 132] and TPINN [138]. In addition, in order to improve the generalizability and transferability of PaHL methods, Jamali et al. [296] replace the time-consuming part of the physics-based numerical models with more efficient data-driven approaches, increasing their computational efficiency and achieving good generalization capability. Kahana et al. [295] explore the geometry transferability properties of HINTS [190].
@@
Speed and Operability. See Fig. 5 (label 2.4). Real-time and operational hydrodynamic modeling (such as flood forecasting) is crucial for supporting emergency responses and reducing risk and damage [319]. Currently, there are few operational systems that forecast flooding at spatial resolutions that can facilitate emergency preparedness and response actions to mitigate flood impacts. Combining the fast inference of ML models with the reliability of physical models, real-time and operability of hydrodynamic modeling is possible. Specifically, a real-time predictive DCGAN [280] is developed for forecasting floods. The last one-time step-ahead forecast from the DCGAN can serve as a new input for the subsequent time step-ahead forecast, which recursively forms a long lead-time forecast. Random forest [297] serves as a surrogate model for real-time flood prediction at a street scale. The surrogate model is trained using hourly water depths simulated by a 1D/2D hydrodynamic model. In addition, this PiML-based method [291] for reducing the computational load of traditional hydrodynamic methods (such as scalable hydrodynamic modeling) can be used for real-time hydrodynamic modeling. Furthermore, a hybrid framework for real-time flood modeling that merges fluid dynamics with probabilistic techniques [298] is developed to address the excessive computational requirements of high-fidelity real-time modeling. Google’s operational flood forecasting system [223], which integrates hybrid hydrological and hydrodynamic models, is established to deliver precise real-time flood alerts to agencies and the public. This forecasting system includes data validation, stage prediction, inundation modeling, and alert dissemination.
@@
Parameterization and Calibration. See Fig. 5 (label 2.5). The behaviors of hydrodynamic models depend heavily on parameters (such as river bed elevation, channel geometry, precipitation, land use land cover, etc.) that need calibration. However, traditional calibration (such as manually-calibrated process-based models) is highly inefficient and results in nonunique solutions [230]. Thus, many PaML-based parameter learning methods are proposed for hydrodynamic modeling. Specifically, a genetic-k-nearest neighbor hybrid algorithm [299] is proposed to mitigate the impractically high computational efforts associated with conventional calibration search techniques, while still achieving high-quality calibration results. A set of statistical-learning-based data assimilation (SLDA) methods [300] is proposed for estimating parameters in hydrodynamic models. In addition, according to the PiML community, PINNs can be used to solve parameter identification and parameter learning (data-physics-driven parameter discovery). For example, PINNs [301] are actively trained to approximate solutions to the Navier-Stokes equations over a parameter space region, where these parameters define physical properties like domain shape and boundary conditions. Furthermore, differentiable parameter learning [230] is proposed to efficiently learn a global mapping between inputs (and optionally responses) and parameters. Differentiable parameter learning contains a parameter estimation module that maps from raw input information to process-based model parameters, as well as a differentiable process-based model. The use of differentiable parameter learning leads to improved performance, enhanced physical coherence, superior generalizability, and reduced computational cost.
@@
Data Generation. See Fig. 5 (label 2.6). Data generation methods are valuable for simulating hydrodynamic processes under specific conditions. For instance, an integrated model [302] combining hydrodynamics and ML is proposed to predict water level dynamics as an indicator of compound flooding risk in a data-limited delta region. In this integrated approach, a hydrodynamic model first simulates various scenarios of compound flooding, and the outputs are then used to train the ML model. A hybrid framework [309] that integrates a hydrodynamic model, rapid flood model, and ML model is used for flood susceptibility assessments. Flood inventory data is generated by the hydrodynamic model. However, traditional methods in hydrodynamics for generating data typically involve running physical simulations or conducting experiments, which are often time-consuming. Hence, there is growing interest in generative ML approaches that can learn data distributions in unsupervised settings, potentially generating novel data beyond what traditional methods can produce. For example, FloodGAN [303] utilizes an image-to-image translation approach where the model learns to generate 2D inundation predictions conditioned on heterogeneous rainfall distributions, employing a minimax game between two adversarial networks. Conditional GAN [304] can also be trained to simulate fluid flow based solely on observations, without knowledge of the underlying governing equations. However, a well-known drawback of GANs is their high sample complexity and lack of diversity. Therefore, several GANs based on PeML and PiML have been proposed to leverage prior knowledge of physics for hydrodynamic processes. For example, GAN-based models for simulating turbulent flows can be enhanced by incorporating conservation laws into the loss function [305]. Physics-informed GAN [306] is employed to learn a functional prior from historical data. In addition, some generative modeling approaches are developed to efficiently model data spaces with known invariances [307, 308].
@@
However, most hydraulic problems use traditional hydrodynamic models to generate data. With the current development of generative models (such as GANs [41], diffusion models [320], and generative pre-trained transformer [321, 322]) in the PaML community, the rapid generation of physical data based on generative models will gradually dominate.
@@
Uncertainty Quantification in Hydrodynamic Modeling. See Fig. 5 (label 2.7). Estimating uncertainty in hydrodynamic modeling is important for many applications. For instance, understanding, characterizing, and quantifying the impact of model and parametric uncertainties is crucial for making informed risk-based decisions in response to intense rainfall events. Monte Carlo, first-order second-moment, and metamodeling are three popular methods to estimate the uncertainty of hydrodynamic models [323]. Abbaszadeh et al. [324] explore multiple origins of uncertainties across distinct layers of hydrometeorological and hydrodynamic model simulations as well as their complex interactions and cascading effects (e.g., uncertainty propagation) in forecasting flooding. Specifically, recent advances in PDgML techniques have shown the benefits of using remote sensing data to correct elevation errors associated with building artifacts, flood defense structures, forests, and wetlands (such as object-based correction [310], FABDEM [311], and random forest regression [312]). In addition, Bayesian estimation/inversion methods are commonly employed to quantify and mitigate modeling uncertainties, such as employing a Markov chain Monte Carlo approach [313] or Bayesian calibration [314]. Furthermore, PiML and PeML approaches have been gaining significant attention in the scientific community, as these methods can alleviate the computational burden and enhance the efficiency required for complex hydrodynamic modeling. For example, the physics-informed Gaussian process regression method [315] combined with the multi-level Monte Carlo for reducing the cost of estimating uncertainty is designed for real-time flood depth forecasting. In addition, a deep probabilistic generative model [316] implements a physics-based loss for uncertainty quantification, ensuring adherence to the structure imposed by PDEs. Bayesian PINNs [118] is proposed by using a Bayesian NN, which naturally encodes uncertainty. Bayesian PINNs [118] incorporate a PDE constraint to enforce the governing laws of the system as a prior for the Bayesian Network, enhancing prediction accuracy in scenarios with significant noise through physics-based regularization. Another direction is a combination of physics knowledge and ML for uncertainty quantification. Hybrid frameworks that combine ML and hydrodynamic models can identify spatio-temporal features from historical flood events, aiding in predicting spatially distributed water depth and flood inundation extent caused by fluvial and coastal drivers [317].
@@
Integrating prior physics knowledge into ML for uncertainty quantification in hydrodynamic modeling has the potential to better characterize uncertainty. Further research in this area is recommended to fully leverage the benefits of the PaML community. PaML models have gained popularity in recent years as an efficient tool for hydrodynamic modeling, designing these emulators and their configurations remains challenging due to the uncertainties inherent in physical models, ML models, and their hybrid methods. There have been insufficient studies to rely on PaML and convincingly demonstrate its utility and effectiveness in solving challenging hydrodynamic problems.
Figure 6: The operational flood forecasting system based on HydroPML.
Table 4: Specific examples of HydroPML-based application highlights.
Applications Examples Model Category Model Details Performances
Rainfall-induced landslide modeling and forecast LandslideCast: A DL method for dynamic process modeling of real landslides [325] PDgML (Deep Operator Networks) [Model] The supervised FNO method is employed to learn the landslide dynamics process.
[Input/framework/output] The model inputs include topography, initial body height, base friction angle, and cohesion. The network maps these inputs to a high-dimensional representation space, iterates and updates them through four Fourier layers, and then maps the results to the output space. The model outputs the time series depth and speed, which reflect the dynamic characteristics of the landslide.
The PDgML method replicates landslide dynamics with high performance, demonstrating excellent computational efficiency, generalization, and zero-shot downscaling capabilities.
Flood modeling and forecast FloodCast: Large-scale flood modeling and forecasting [35] PiML (PINOs) [Model] A geometry-adaptive physics-informed neural solver (GeoPINS) is introduced, which benefits from not requiring training data in PINNs and features a fast, accurate, and resolution-invariant architecture using FNO.
[Input/framework/output] GeoPINS uses input channels composed of coordinates, time domain, and initial conditions of flood height, which are first lifted to a higher-dimensional representation by an MLP. Several Fourier layers then extract efficient spatiotemporal representations, and the outputs (flood depth) are obtained via MLP projection. GeoPINS is trained with a geometry-adaptive physics-constrained loss, which includes physics loss for controlling the residuals of 2-D depth-averaged shallow water equations and data loss for measuring the mismatch between the NN prediction and the initial/boundary conditions.
The experimental results for the 2022 Pakistan flood demonstrate that the GeoPINS enables high-precision, large-scale flood modeling. For instance, traditional hydrodynamics and GeoPINS exhibit exceptional agreement during high water levels.
Rainfall-runoff modeling and forecast MC-LSTM: Mass-conserving LSTM for rainfall-runoff modeling [166, 241] PeML (PPeML) [Model] MC-LSTM extends the inductive bias of the traditional LSTM to uphold conservation laws, ensuring mass input is conserved through modifications to the recurrent structure.
[Input/framework/output] The input of MC-LSTM includes precipitation, with auxiliary inputs comprising daily minimum and maximum temperature, solar radiation, vapor pressure, and 27 basin characteristics related to geology, vegetation, and climate. MC-LSTM directly learns the relationship between these inputs and streamflow or other output fluxes.
MC-LSTM exhibits superior performance compared to other mass-conserving hydrological models, including the mesoscale hydrologic model (mHM), VIC, HBV, and others.
Hydrodynamic process understanding Physics-aware downsampling method: A DL method for scalable flood modeling [194, 326] PaHL (Neural–numerical) [Model] Physics-aware downsampling method for scalable 2D hydrodynamic modeling is proposed by minimizing the distance between the solutions on the fine and coarse grids.
[Input/framework/output] The model takes fine grid terrain as input and employs a downsampling neural network to transform it into coarse grid terrain. Subsequently, a parallel numerical solver with initial and boundary conditions on fine grid terrain is used to provide a supervised signal for the predictions on the coarse grids. This approach ensures that flood predictions on the coarse grids align closely with those on the fine grid.
For hydrodynamic modeling, significant reductions in computational cost are achieved without compromising solution accuracy.
@@
Rainfall-runoff-inundation modeling and forecast The operational flood forecasting system based on HydroPML Comprehensive Four main modeling components include: (1) data management based on parameterization and calibration, (2) a rainfall-runoff forecasting model based on PaML, (3) a flood inundation model based on hydrodynamic models, and (4) alerts to the affected population. Real-time flood forecast and flood level information are provided. The alerts and warnings comprise information about the forecasted water level, the inundation map, and the inundation depth.
3.3 Proposed HydroPML-based Application Highlights
To illustrate the practical utility and performance of the HydroPML platform, we provide several specific examples of HydroPML-based application highlights, as shown in Table 4. These include rainfall-induced landslide modeling and forecasting, flood modeling and forecasting, rainfall-runoff modeling and forecasting, hydrodynamic process understanding, and rainfall-runoff-inundation modeling and forecasting. For rainfall-induced landslide modeling and forecasting, a supervised deep operator network (PDgML) [325] is employed in HydroPML to learn the landslide dynamics process, replicating landslide dynamics with high performance. For flood modeling and forecasting, GeoPINS (PiML) [35] is used in HydroPML for large-scale flood forecasting. The experimental results for the 2022 Pakistan flood demonstrate that GeoPINS enables high-precision, large-scale flood modeling. For rainfall-runoff modeling and forecasting, MC-LSTM (PeML) [166, 241] is employed in HydroPML for rainfall-runoff modeling, exhibiting superior performance compared to other mass-conserving hydrological models. For hydrodynamic process understanding, the physics-aware downsampling method (PaHL) [194, 326] is used in HydroPML for scalable 2D hydrodynamic modeling, achieving significant reductions in computational cost without compromising solution accuracy. The details of these models (such as category, input, framework, and output) are introduced in detail in Table 4. The following presents a detailed application case of PaML’s comprehensive methods: rainfall-runoff-inundation modeling and forecast.
Based on the HydroPML technology, an operational flood forecasting (rainfall-runoff-inundation) system is developed. As shown in Fig. 6, four key modeling components are integrated into flood warning systems: (1) data management based on parameterization and calibration, (2) a rainfall-runoff forecasting model based on PaML, (3) a flood inundation model based on hydrodynamic models, and (4) alerts to the affected population.